# FedVSR: Towards Model-Agnostic Federated Learning in Video Super-Resolution

### v2

@@ 5 Experiments
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
VRT [30] FedAvg[36] 29.71 0.8522 0.1916 24.88 0.7680 0.2446 28.85 0.8282 0.2200 24.60 0.7516 0.2557
FedProx[27] 28.82 0.8281 0.2133 24.18 0.7341 0.2668 28.64 0.8236 0.2222 24.10 0.7256 0.2713
SCAFFOLD[21] 28.59 0.8181 0.2228 23.99 0.7182 0.2737 28.17 0.8074 0.2318 24.04 0.7208 0.2739
FedVSR (Ours) 29.86 0.8562 0.1908 24.92 0.7682 0.2393 29.03 0.8339 0.2123 24.73 0.7584 0.2492
RVRT [29] FedAvg[36] 31.30 0.8877 0.1536 25.96 0.8156 0.1974 29.89 0.8578 0.1867 25.67 0.8040 0.2065
FedProx[27] 30.09 0.8618 0.1899 25.20 0.7775 0.2330 29.59 0.8526 0.1943 25.02 0.7730 0.2353
SCAFFOLD[21] 29.63 0.7967 0.2101 24.79 0.7317 0.2407 26.71 0.7142 0.2590 23.08 0.6441 0.3018
FedVSR (Ours) 31.54 0.8922 0.1455 26.05 0.8191 0.1889 30.90 0.8811 0.1576 25.90 0.8131 0.1910
IART [54] FedAvg[36] 30.56 0.8729 0.1695 25.43 0.7949 0.2118 29.52 0.8493 0.1909 25.04 0.7774 0.2272
FedProx[27] 29.65 0.8500 0.1683 24.72 0.7599 0.2307 29.30 0.8444 0.1958 24.42 0.7505 0.2293
SCAFFOLD[21] 30.29 0.8678 0.1692 25.18 0.7857 0.2084 29.26 0.8413 0.1992 24.85 0.7690 0.2278
FedVSR (Ours) 30.88 0.8812 0.1556 25.54 0.8012 0.1990 29.76 0.8536 0.1887 25.26 0.7850 0.2220
Table 1: PSNR ($\uparrow$), SSIM ($\uparrow$), and LPIPS ($\downarrow$) comparison of FL algorithms across different VSR models and datasets under varying levels of heterogeneity.
5.1 setup
Dataset:
Following [30, 29, 54], one of the most commonly used datasets for training VSR is REDS [40]. It contains 240, 30, and 30 sequences for training, validation, and testing, respectively, with each sequence consisting of 100 frames at a resolution of 720 × 1280. The corresponding low-resolution (LR) frames are generated through bicubic downsampling from the high-resolution (HR) images. This dataset features a variety of natural and handmade objects, locations, and diverse content, making it heterogeneous in decentralized settings [40]. To further evaluate FedVSR and assess its robustness in more realistic, highly heterogeneous scenarios, we conducted additional experiments using Kinetics-400 [22], a dataset primarily used for human action recognition. It includes 400 action classes with 306,245 videos, each approximately 10 seconds long, sourced from unique YouTube videos. Since Kinetics-400 contains videos of varying resolutions, we selected only those with a 720 × 1280 resolution to ensure a fair comparison with REDS. From the 400 classes, we randomly chose 40, and for each class, selected six videos, extracting 100-frame clips to match REDS’ format. Kinetics-400 is particularly well-suited for testing FedVSR under more extreme heterogeneity because, unlike professionally filmed datasets, its videos often exhibit camera shake, lighting variations, shadows, and background clutter [22]. This makes it a challenging but realistic benchmark for assessing the model’s adaptability in diverse and less controlled environments. For consistency with our experimental setup, we refer to REDS as the ”less heterogeneous” dataset and Kinetics-400 as the ”more heterogeneous” dataset, given the increased variability in content, motion, and environmental conditions in the latter.
Implementation Details:
To maintain our objective of developing a model-agnostic framework, we did not alter the official parameters of the baseline VSR models. Instead, we used their recommended settings for training. We created 40 clients, each assigned six videos in both experimental setups. In the more heterogeneous setup, each client received videos from a single class; for example, one client had only ”archery” videos, while another had only ”making pizza” videos. In contrast, in the less heterogeneous setup, the six videos were randomly assigned to each client. We set the total number of global rounds to 100, with each client performing one local round (epoch) per global round. In each round, we randomly selected four clients to perform training. After completing their local training, these clients sent their updated weights to the server for aggregation. All experiments were conducted using bi-cubic (BI) degradation.
Baselines:
For our baseline, we use a combination of FedAvg[36], FedProx[27], and SCAFFOLD[21] from the FL side, along with VRT[30], RVRT[29], and IART[54] from VSR. We selected these FL methods as they are widely accepted and frequently used to benchmark core algorithmic performance in recent literature. While newer approaches like MOON[24], FedDyn[1], FedDisco[57], FedLAW[28] exist, they are rely on task-specific assumptions such as logits, class-wise statistics, or server-side data proxies that do not translate well to low-level, regression-based tasks like VSR. On the other hand, exisitng FL-SR models such as FedSR, p²FedSR, and FedMRI are not directly comparable to FedVSR for two key reasons. First, they are tightly coupled to a specific model architecture: they require dividing the model into predefined blocks (“shallow–deep–upsampler”), which limits flexibility and demands structural changes to existing SR models. Second, these methods are stateful meaning each client must store additional data across training rounds, such as degradation descriptors, deep feature extractors, or previous model weights. This leads to significantly higher storage, computation, and coordination overhead. In contrast, FedVSR is designed to be both model-agnostic and stateless, making it simpler, more lightweight, and easier to apply in diverse real-world scenarios.
Evaluation Metrics:
The evaluation metrics used for the experiments are Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) and Learned Perceptual Image Patch Similarity (LPIPS).
Figure 2: PSNR values across different rounds for various test sets under different settings for VRT [30], RVRT [29], and IART [54].
Reproducibility:
To ensure our experiments are reproducible, we fixed the random seed, which doesn’t affect client selection, model initialization, etc. We used four NVIDIA A6000 GPUs for training, with each model under a federated learning algorithm taking an average 46 hours to train. For the baselines alone, this consumed 36 days (2 experiments × 9 models+algorithms × 2 days per run).
5.2 Results and Comparison
Quantitative Results:
As shown in Table 1, our proposed FedVSR algorithm consistently outperforms other federated learning (FL) algorithms across all tested models and datasets. On average, FedVSR surpasses existing FL methods by a margin of 0.61 dB in PSNR and 0.022 in SSIM and 0.022 in LPIPS in less heterogeneous settings, while in more heterogeneous settings, the margins are 0.66 dB in PSNR and 0.028 in SSIM and 0.019 in LPIPS, demonstrating its robustness under challenging conditions. Among the evaluated models, RVRT benefits the most from FedVSR, achieving the highest performance margin. In contrast, IART shows the lowest overall improvement but still maintains a competitive edge over the baseline FL methods.
The observed performance trends in our experiments align with existing findings on FedProx and SCAFFOLD, further validating our results. As shown in Fig.2, FedVSR consistently achieves superior PSNR values across different VSR models and datasets, outperforming FedAvg, FedProx, and SCAFFOLD, particularly in later training rounds. This aligns with prior research on FedProx, which highlights its failure to reach the stationary points of the global optimization objective, even under homogeneous data distributions. The study suggests that FedProx’s limiting points may remain far from the true minimizer of the global empirical risk [48], which could explain its suboptimal performance in our experiments. Although FedProx may still generalize well by aggregating local data effectively in certain conditions, its weaker convergence properties become apparent in more heterogeneous scenarios. We also observe that FedProx is highly sensitive to the $\mu$ value during training. The original paper suggests $\{0.001,0.01,0.1,1\}$ as candidate values and finds that $\mu=1$ performs better in most cases. However, we notice that for any $\mu$ greater than 0.001, the VSR model struggles significantly with learning. This suggests that the relationship is more complex and requires further investigation and analysis.
Similarly, our findings regarding SCAFFOLD’s inefficacy in highly heterogeneous settings corroborate previous research that highlights its susceptibility to instability under heterogeneous data [4]. The significant performance drops and stagnation observed in our experiments, particularly with the RVRT model, align with the reported issues of SCAFFOLD transmitting nearly twice the data of other algorithms due to its reliance on control variates. This added communication burden does not necessarily translate to better convergence, as our results show that SCAFFOLD struggles to optimize effectively in diverse settings. Prior work [4] also notes that SCAFFOLD is highly prone to sudden performance breakdowns when gradient clipping is not employed, which is reflected in our results where its PSNR fluctuates or plateaus early in highly heterogeneous scenarios. The increased variance in SCAFFOLD’s performance under skewed data distributions further supports our findings that it fails to adapt smoothly across different client datasets, reinforcing its limitations in federated VSR tasks.
Qualitative Results:
As showen in Fig. 1 clearly demonstrate the limitations of FedAvg, FedProx, and SCAFFOLD in accurately reconstructing fine details. While these methods reduce some pixelation from the LQ input, they all fail to recover the textures and sharp edges present in the ground-truth. FedAvg produces overly smooth results, blurring important structural details such as the wooden lattice and frame edges. FedProx exhibits similar blurring, struggling to restore high-frequency textures. SCAFFOLD, not only fails to produce a sharp reconstruction but also introduces noticeable artifacts, distorting the texture of the wooden lattice and making the output appear unnatural. In contrast, FedVSR delivers significantly sharper and more realistic reconstructions, clearly outperforming the other approaches. Due to space limitations, additional qualitative and quantitative results are provided in the supplementary material 8.
@@ 5.3 Ablation Study
Evaluating $\mathcal{L}_{\text{HiFr}}$ Loss and Adaptive Aggregation
To further analyze the impact of the DWT loss term and loss-aware adaptive aggregation, we conducted a series of experiments. Using FedAvg as a baseline, we added each component separately to evaluate their individual effects. To assess the impact of the $\mathcal{L}_{\text{HiFr}}$ term, we applied it to FedAvg and observed the changes (F-$\mathcal{L}_{\text{HiFr}}$). Similarly, to evaluate loss-aware adaptive aggregation, we first trained VSR models using a greedy aggregation strategy (F-GLA) that prioritizes low-loss updates in all rounds. We then compared this approach to an adaptive version that incorporates a decay mechanism (F-ALA). As shown in Table 2, the DWT loss term consistently outperforms loss-aware adaptive aggregation in all cases. Additionally, both methods individually surpass FedAvg [36]. However, in some cases, the greedy aggregation strategy resulted in only marginal improvements or even a drop in overall performance especially in ”more heterogeneous” setting based on PSNR and SSIM metrics. These findings further highlight the importance of high-frequency information not only for improving overall performance but also as a reliable metric for identifying valuable client weights. Due to space limitations, LPIPS metric results are provided in the supplementary material 7
@@
Table 2: Ablation study on the impact of $\mathcal{L}_{\text{HiFr}}$ and adaptive aggregation on PSNR and SSIM under different heterogeneity settings.
Optimal $\lambda$s
Using a grid search, we explored different values for $\lambda_{\text{VSR}}$ and $\lambda_{\text{HiFr}}$ to determine the optimal ones. We varied each parameter from 1 to 0.1 in steps of 0.1 under both bounded and unbounded conditions. Due to the high computational cost of the experiments, we trained the VSR models for only 20 rounds and selected the best-performing configurations. Our results show that for VRT[30], the optimal values for both parameters are 1 For RVRT[29], the best values are 1 and 0.5, while for IART[54], the optimal values are 0.5 and 0.5, respectively.
Empirical Fisher for Estimating Weights Importance
Inspired by [35], we explored estimating the importance of each weight in client models to enhance aggregation using Empirical Fisher (EF) Information. EF approximates the Fisher information matrix by leveraging gradients collected during backpropagation after training. While this method led to slight performance improvements, it proved computationally too expensive for clients, making it impractical for our setting. A more efficient variant of EF was introduced in [52]; however, as it is tailored for classification tasks, we leave its adaptation to VSR for future work.
@@ 6 Conclusion
In this work, we introduced FedVSR, the first FL framework specifically designed for VSR, addressing privacy concerns while preserving high-quality performance. Unlike existing FL approaches that struggle with low-level vision tasks, our model-agnostic and stateless design ensures adaptability across various VSR architectures without prior model knowledge or historical weight storage. By incorporating a lightweight loss term, we effectively guide local updates and enhance global aggregation, resulting in notable performance improvements over general FL baselines. Our method functions as a plug-and-play solution, and extensive experiments confirm its superiority over existing FL approaches, demonstrating its potential for real-world, privacy-sensitive applications. Furthermore, our study highlights key challenges in applying FL to low-level vision tasks, offering valuable insights for future research in this under-explored domain. As the first FL-based approach for VSR, this work establishes a strong baseline and provides a clear research path for future advancements. Despite promising results, a gap remains between FL-based VSR and centralized approaches. Future research could refine the aggregation strategy with a more intelligent, task-specific approach tailored for low-level vision. Additionally, our loss-guided optimization could be further enhanced to improve stability and better address existing challenges in FL-based VSR.
Supplementary Material
7 Extended Evaluation of HiFr Loss and Adaptive Aggregation Using LPIPS
To complement the analysis based on PSNR and SSIM discussed earlier in the paper, we present here an additional evaluation using LPIPS, which serves as a perceptual similarity metric. This acts as an extension to the previously discussed metrics and provides a more complete picture of model performance. Due to space limitations, we include this analysis here.
To evaluate the impact of the DWT loss term and the loss-aware adaptive aggregation strategies, we conducted a set of ablation experiments using FedAvg as the baseline. Each component was added independently to isolate its individual effect. Specifically, we assessed the contribution of the $\mathcal{L}_{\text{HiFr}}$ term by applying it to FedAvg (F-$\mathcal{L}_{\text{HiFr}}$), and examined two aggregation variants: a greedy loss-aware approach (F-GLA) that prioritizes updates from clients with lower losses, and an adaptive version (F-ALA) that incorporates a decay factor to regulate client influence over time.
As shown in Table 3 with LPIPS results, the observed trends remain largely consistent across metrics. The DWT loss term consistently provides more significant gains compared to the aggregation strategies, and both methods individually outperform the standard FedAvg baseline. However, similar to PSNR and SSIM, the greedy strategy sometimes leads to only marginal improvements or even minor performance degradation, especially under high heterogeneity. These findings underscore the importance of high-frequency information not only for boosting perceptual quality but also for reliably identifying valuable client updates.
Model Setting Less Heterogeneous More Heterogeneous
REDS Vid4 REDS Vid4
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
VRT F-$\mathcal{L}_{\text{HiFr}}$ 29.80 0.8543 0.1910 24.91 0.7681 0.2399 28.98 0.8322 0.2130 24.70 0.7577 0.2500
F-ALA 29.75 0.8531 0.1912 24.89 0.7681 0.2411 28.90 0.8313 0.2140 24.64 0.7559 0.2520
F-GLA 29.70 0.8523 0.1914 24.86 0.7670 0.2436 28.83 0.8279 0.2198 24.55 0.7500 0.2555
RVRT F-$\mathcal{L}_{\text{HiFr}}$ 31.42 0.8888 0.1506 26.02 0.8178 0.1910 30.24 0.8617 0.1612 25.84 0.8042 0.1978
F-ALA 31.38 0.8880 0.1518 26.00 0.8171 0.1931 30.19 0.8601 0.1728 25.79 0.8041 0.1992
F-GLA 31.31 0.8879 0.1527 25.95 0.8155 0.1948 29.88 0.8576 0.1807 25.60 0.8032 0.2011
IART F-$\mathcal{L}_{\text{HiFr}}$ 30.79 0.8801 0.1588 25.46 0.8001 0.1999 29.64 0.8508 0.1899 25.21 0.7837 0.2231
F-ALA 30.70 0.8789 0.1591 25.44 0.7998 0.2101 29.60 0.8500 0.1903 25.18 0.7831 0.2246
F-GLA 30.57 0.8730 0.1690 25.42 0.7951 0.2119 29.50 0.8482 0.1929 24.99 0.7761 0.2284
Table 3: Ablation study on the impact of $\mathcal{L}_{\text{HiFr}}$ and adaptive aggregation on PSNR, SSIM, and predicted LPIPS under different heterogeneity settings.
8 Further Analysis: Performance Metrics and Visual Comparisons
This section provides additional performance evaluations for our proposed method. Figures 3, 5, and 7 illustrate the PSNR, SSIM and LPIPS trends across different rounds for multiple test sets under various configurations of VRT [30], RVRT [29], and IART [54]. These results further highlight the effectiveness of our approach in enhancing the performance across different models.
Figures 4, 6, and 8 provide qualitative visual comparisons on various test frames. Each figure presents a reference ground truth (GT) frame alongside low-quality (LQ) input and the outputs generated by different Federated Learning strategies, including FedAvg, FedProx, SCAFFOLD, and our method (FedVSR). The reconstructions show that FedVSR produces clearer and more visually faithful outputs, particularly in preserving finer details and reducing artifacts.
Figure 3: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for VRT.
Refer to caption (a) Frame 000, Clip 015, REDS
Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 32.41, 0.8981, 0.1456 Refer to caption (d) FedProx[27] 32.10, 0.8935, 0.1511
Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 28.85, 0.7625, 0.1934 Refer to caption (g) FedVSR (Ours) 33.79, 0.9169, 0.1151
Figure 4: Comparison of RVRT output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.
Figure 5: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for RVRT.
Refer to caption (a) Frame 040, Clip 1, Vid4
Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 21.15, 0.7185, 0.3520 Refer to caption (d) FedProx[27] 20.43, 0.6735, 0.4101
Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 20.41, 0.6679, 0.3803 Refer to caption (g) FedVSR (Ours) 21.19, 0.7200, 0.3433
Figure 6: Comparison of VRT output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.
Figure 7: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for IART.
Refer to caption (a) Frame 070, Clip 000, REDS
Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 27.87, 0.8211, 0.2228 Refer to caption (d) FedProx[27] 27.54, 0.8132, 0.2247
Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 27.69, 0.8151, 0.2207 Refer to caption (g) FedVSR (Ours) 28.07, 0.8319, 0.2014
Figure 8: Comparison of IART output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.

### v3

@@ 5.3.1 Additional Baseline: FedSR vs FedVSR
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
FedSR [50] 29.70 0.8518 0.1939 24.89 0.7676 0.2470 28.85 0.8297 0.2156 24.70 0.7611 0.2571
FedVSR 29.96 0.8595 0.1807 24.97 0.7745 0.2284 29.12 0.8371 0.2010 24.78 0.7646 0.2378
Table 3: Quantitative comparison of FedSR [50], and FedVSR under different heterogeneity settings. Metrics include PSNR, SSIM, and LPIPS for REDS and VID4 datasets.
5.3.2 Robustness to Communication Failures
To evaluate the robustness of our method under unreliable communication, we conduct a communication reliability test by simulating scenarios in which a subset of clients fail to upload their updated weights to the server. Specifically, we consider failure rates of 25%, 50%, and 75%, meaning that in each round, the corresponding fraction of clients’ updates are discarded. We test this setting using two models FedVSR, and FedAvg. This setup allows us to compare how each approach handles upload failures and missing updates, and how well they maintain performance under increasingly adverse communication conditions. Fig.  4 illustrates the resulting performance for each method across the different failure rates. Our results indicate that FedVSR demonstrates greater resilience, maintaining performance even at high failure rates, whereas FedAvg shows significant degradation as the failure rate increases. These findings highlight the advantage of our approach in real-world scenarios where client-server communication may be unreliable.
Figure 4: FedVSR vs. FedAvg under client upload failures (0–75%), showing FedVSR’s higher robustness.
5.3.3 Client Population Stress Test
To evaluate the scalability of our method, we conduct a client population stress test in which we increase the number of clients by three and six times compared to the main experiment 5.1.2, while keeping the client participation rate constant (10%). We compare the performance of FedVSR and FedAvg under these conditions. This experimental setup allows us to investigate whether our method can maintain high performance when faced with a substantially larger client population, without requiring changes to participation strategies. Fig. 5 shows the results of this test, indicating that FedVSR consistently outperforms FedAvg even as the number of clients increases, demonstrating the scalability and robustness of our approach in large-scale federated learning scenarios.
Figure 5: FedVSR vs. FedAvg under client population stress test.
TC PR Algorithm REDS [35] Vid4 [26]
PSNR SSIM LPIPS PSNR SSIM LPIPS
Less Heterogeneous
120 3.3% FedAvg 29.21 0.8261 0.2136 24.36 0.7362 0.2716
FedVSR 29.36 0.8378 0.1920 24.47 0.7472 0.2431
240 1.7% FedAvg 28.23 0.8027 0.2327 23.94 0.7079 0.2893
FedVSR 28.59 0.8155 0.2093 24.10 0.7225 0.2562
More Heterogeneous
120 3.3% FedAvg 28.36 0.8028 0.2453 24.09 0.7205 0.2839
FedVSR 28.54 0.8160 0.2136 24.28 0.7376 0.2531
240 1.7% FedAvg 26.95 0.7562 0.2979 23.18 0.6641 0.3358
FedVSR 27.23 0.7742 0.2474 23.44 0.6881 0.2839
Table 4: Extreme test results for FedAvg and FedVSR under different heterogeneity settings. Here, TC denotes the total number of clients and PR denotes the participation rate.
5.3.4 Extreme Case: Large Client Population with Low Participation
To investigate the behavior of our method under extreme conditions, we conduct a test combining a substantially increased client population with very low participation rates. Specifically, we evaluate scenarios with three and six times more clients than the main experiment 5.1.2, while reducing the participation rate to approximately one-third (3.3%) and less than one-fifth (1.7%) of the original rate (baseline 10%). We compare the performance of FedVSR and FedAvg under these conditions to assess how well each method handles extreme sparsity in client updates combined with a large-scale client population. Table 4 presents the results, demonstrating that FedVSR maintains superior performance and robustness even in these challenging scenarios, whereas FedAvg experiences significant performance degradation. These findings highlight the resilience and scalability of FedVSR under highly adverse federated learning conditions.
5.3.5 Effect of Local Epochs on Communication-Efficient Performance
To investigate the trade-off between local computation and communication efficiency, we compare the performance of FedVSR and FedAvg after 20 global rounds using different numbers of local epochs per client. These experiments are measured relative to our main experiment, which consists of 100 global rounds with 1 local epoch per client. The results are presented in the Fig.  6 as percentages of the main experiment’s performance. This comparison allows us to assess whether increasing the number of local epochs can compensate for fewer global rounds while highlighting the differences between FedVSR and FedAvg under limited communication.
Figure 6: Effect of local epochs on FedVSR and FedAvg (% of 100-round FedVSR with 1 local epoch).
5.3.6 Resource and Efficiency Overhead
Table 5 reports the computation and memory overhead of different FL algorithms relative to FedAvg. We select FedAvg as the baseline because it is a vanilla FL algorithm that introduces no additional overhead. To ensure accuracy, each measurement was repeated multiple times. We also report results as relative percentages, since our experiments involved three different VSR models with varying computational requirements. A value of 0% indicates no overhead beyond FedAvg, while higher values indicate additional computational or memory costs. Presenting the results in this way provides a clearer comparison across models. For these measurements, we assume an ideal network, as the communication cost of transmitting parameters is negligible compared to training time, as discussed earlier 5.1.5. If we were to consider communication cost, in FedVSR each client only needs to send a single scalar, unlike SCAFFOLD or FedSR, which incur a much heavier overhead. As shown, FedVSR not only delivers strong performance but also incurs minimal overhead compared to its counterparts.
   Model    Phase    Latency    Energy    GPU Mem    GPU Util
   FedVSR    Training    0.25    0.73    0.01    0.17
   Aggregation    5.39    4.36
   SCAFFOLD    Training    10.56    11.35    5.53    4.85
   Aggregation    17.44    21.12
   FedProx    Training    17.32    19.82    0.71    2.93
   Aggregation    39.74    39.20
   FedSR    Training    1.90    4.61    3.16    11.92
   Aggregation    314.62    620.42
   FedMedian    Training    0    0    0    0
   Aggregation    2.15    1.51
Table 5: Computation and memory overhead of different FL methods relative to FedAvg (%).GPU memory and utilization are reported for both Training and Aggregation phases.
@@ 5.4 Ablation Study
5.4.1 Evaluating High Frequency Loss on Centralized Setting
The results clearly demonstrate that adding $\mathcal{L}_{\text{HiFr}}$ to centralized training negatively impacts performance across all metrics. This degradation is observed consistently in both the less and more heterogeneous data splits. This outcome confirms that the $\mathcal{L}_{\text{HiFr}}$ loss is not a generic VSR improvement. Rather, it is a necessary intervention specifically designed to counteract the core failings of Federated Learning in low-level vision tasks. In centralized training, the model has sufficient, high-bandwidth data access to efficiently capture high-frequency details using the standard VSR loss. Introducing the DWT-based loss in this setting acts as a redundant constraint, likely leading to over-regularization or gradient noise, which harms the final model quality. Our finding establishes that $\mathcal{L}_{\text{HiFr}}$ is a crucial, non-trivial component of the FedVSR framework designed exclusively to fill the performance gap created by the inherent constraints of federated learning for VSR.
   Algorithm    REDS [35]    Vid4 [26]
   PSNR    SSIM    LPIPS    PSNR    SSIM    LPIPS
   Less Heterogeneous
   VRT    32.20    0.9007    0.1342    26.31    0.8266    0.1780
[…]
   More Heterogeneous
   VRT    31.06    0.8688    0.1505    25.25    0.7900    0.1993
[…]
Table 6: Ablation study comparing VRT performance with and without the $\mathcal{L}_{\text{HiFr}}$ loss under centralized training.
5.4.2 Evaluating High Frequency Loss and Adaptive Aggregation
To further analyze the impact of the DWT loss term and loss-aware adaptive aggregation, we conducted a series of experiments. Using FedAvg as a baseline, we added each component separately to evaluate their individual effects. To assess the impact of the $\mathcal{L}_{\text{HiFr}}$ term, we applied it to FedAvg and observed the changes (F-$\mathcal{L}_{\text{HiFr}}$). Similarly, to evaluate loss-aware adaptive aggregation, we first trained VSR models using a greedy aggregation strategy (F-GLA) that prioritizes low-loss updates in all rounds. We then compared this approach to our adaptive aggregation (F-ALA). As shown in Table 7, the DWT loss term consistently outperforms loss-aware adaptive aggregation in all cases. Additionally, both methods individually surpass FedAvg. However, in some cases, the greedy aggregation strategy resulted in only marginal improvements or even a drop in overall performance especially in ”more heterogeneous” setting based on PSNR and SSIM metrics. These findings further highlight the importance of high-frequency information not only for improving overall performance but also as a reliable metric for identifying valuable client weights.
@@ 5.4.2 Evaluating High Frequency Loss and Adaptive Aggregation
Table 7: Ablation study on the impact of $\mathcal{L}_{\text{HiFr}}$ and adaptive aggregation under different heterogeneity settings.
@@ 6 Conclusion
In this work, we introduced FedVSR, the first FL framework specifically designed for VSR, addressing privacy concerns while preserving high-quality performance. Unlike existing FL approaches that struggle with low-level vision tasks, our model-agnostic and stateless design ensures adaptability across various VSR architectures without prior model knowledge or historical weight storage. By incorporating a lightweight loss term, we effectively guide local updates and enhance global aggregation, resulting in notable performance improvements over general FL baselines. Our method functions as a plug-and-play solution, and extensive experiments confirm its superiority over existing FL approaches, demonstrating its potential for real-world, privacy-sensitive applications. Furthermore, our study highlights key challenges in applying FL to low-level vision tasks, offering valuable insights for future research in this under-explored domain. As the first FL-based approach for VSR, this work establishes a strong baseline and provides a clear research path for future advancements. Despite promising results, a gap remains between FL-based VSR and centralized approaches. Future work could improve aggregation with task-specific strategies for low-level vision and enhance our loss-guided optimization for greater stability and effectiveness in FL-based VSR.
Future Work
A key direction for future work is extending FedVSR to handle higher-resolution videos and scenarios where clients may contribute data at differing native resolutions.